John invested the following amounts in three stocks: Security Investment Beta Stock A $134,061 2.28 Stock B $856,728 1.90 Stock C $735,080 0.94 Calculate the beta portfolio.

Round the answers to two decimal places.
Answer #1

Total value = 134,061 + 856,728 + 735,080 = 1,725,869

Beta portfolio = (134,061 / 1,725,869)*2.28 + (856,728 / 1,725,869)*1.9 + (735,080 / 1,725,869)*0.94

Beta portfolio = 0.177104 + 0.943167 + 0.400364

Beta portfolio = 1.52